### 19 Mar 2026 v2

# An Auditable AI Agent Loop for Empirical Economics: A Case Study in Forecast Combination

Abstract
AI coding agents make empirical specification search fast and cheap, but they also widen hidden researcher degrees of freedom. Building on an open-source agent-loop architecture, this paper adapts that framework to an empirical economics workflow and adds a post-search holdout evaluation. In a forecast-combination illustration, multiple independent agent runs outperform standard benchmarks in the original rolling evaluation, but not all continue to do so on a post-search holdout. Logged search and holdout evaluation together make adaptive specification search more transparent and help distinguish robust improvements from sample-specific discoveries.
Keywords: Agent loops, Autoresearch, Specification search, Researcher degrees of freedom, Responsible AI, Forecast combination
JEL Codes: C53, C52, C18
1 Introduction
Empirical researchers have always explored alternative specifications. AI makes that process cheap and fast. An AI coding agent can rewrite an estimation script, run a fixed evaluator, observe the score, and try again hundreds of times in one session. Once that loop is in place, the practical boundary between coding assistance and undisclosed specification search becomes thin. For a field long concerned with researcher degrees of freedom (Leamer1983; White2000; GelmanLoken2013; Miguel2021), the central question is therefore not whether AI agents can improve an empirical score, but whether their search can be made transparent and auditable.
This paper studies one response. Building on Karpathy2026’s (Karpathy2026) open-source autoresearch, I adapt a simple agentic coding loop to a research protocol for empirical economics. The protocol has four ingredients: a written instruction contract, an immutable evaluator, a single editable script, and a complete experiment log. I then add an explicit post-search holdout evaluation. Together, they do not eliminate adaptive search, but they make it inspectable, bounded, and easier to disclose, while also making validation beyond the search sample explicit. The point is therefore not a new agent architecture, but a way to use an existing one in empirical work with clearer external validation.
The empirical illustration uses the forecast-combination problem of DieboldShin2019. An AI coding agent is asked to find a look-ahead-free real-time tuning rule for peLASSO while leaving the evaluator fixed. On the original rolling evaluation, three independent runs improve on the simple average and on the paper’s ex post peLASSO benchmark. A post-search holdout for 2017Q1–2025Q4 then provides the key validation: two discovered methods continue to outperform standard benchmarks, while one does not.
That split result is the main message. The experiment logs show that nominal hyperparameter tuning readily drifts into broader method discovery. Some departures survive external evaluation; others do not. The value of the framework is therefore twofold: the underlying loop records how the agent searched, and the added holdout evaluation allows the resulting departures to be judged rather than hidden. For empirical economics, that is the relevant promise of agentic search: not automation without discretion, but automation with an audit trail and explicit post-search holdout evaluation.
Recent work on AI for science spans a wide range of autonomy. On one end, systems such as AlphaEvolve (NovikovEtAl2025) and DS-STAR (ChenEtAl2025) use large language models to iteratively rewrite code and refine analyses against automated evaluators. On the other end, systems such as The AI Scientist (LuEtAl2024) and AI co-scientist (GottweisEtAl2025) pursue broader research automation across idea generation, experiment design, analysis, and communication. In economics, Korinek2025 surveys AI agents for research, while DawidEtAl2025 develop broader agentic workflows. Our contribution is narrower: we study evaluator-locked local search within a fixed empirical workflow, where the researcher sets the question, data, evaluator, editable surface, and search budget. The goal is not full automation, but an auditable protocol for standard empirical practice.
The remainder of the paper describes the protocol (Section 2), presents the empirical illustration (Section 3), and concludes (Section 4).
2 An auditable agent-loop protocol
Consider a researcher who starts with a baseline empirical model and wants to explore nearby alternatives using an AI coding agent. [^1] By AI coding agent we mean a general-purpose coding assistant such as Claude Code or OpenAI Codex that plans, writes code, and executes tool calls within a development environment. We formalize the search as follows. Let $C$ denote a written instruction contract that specifies the research objective, admissible modifications, and search budget. Let $\mathcal{T}(C)$ denote the class of admissible editable scripts under $C$. Let the available data be partitioned as $D=(D^{S},D^{H})$, where $D^{S}$ is the search sample and $D^{H}$ is a post-search holdout reserved for external validation. Let $S(\tau;D^{S})$ denote the scalar score returned by an immutable evaluator when script $\tau\in\mathcal{T}(C)$ is run on search sample $D^{S}$. Both $\mathcal{T}(C)$ and $S$ are fixed ex ante by the researcher as part of the search design. In practice, the contract fixes the evaluator, editable surface, and no-look-ahead constraint, but it need not impose a machine-checkable restriction that candidate scripts belong to a particular model family; the effective admissible set $\mathcal{T}(C)$ may therefore be broader than the researcher’s semantic intent. A complete experiment log $L$ records every attempted modification and its outcome. These four elements—contract, evaluator, editable script, and log—constitute the protocol.
When lower scores are preferred, the search problem is
$\tau^{\star}\in\arg\min_{\tau\in\mathcal{T}(C)}S(\tau;D^{S}).$
In practice, the agent sequentially generates and evaluates a finite sequence of candidate scripts $\{\tau_{0},\tau_{1},\ldots,\tau_{K}\}$, each informed by the outcomes of previous attempts, and returns
$\hat{\tau}_{K}\in\arg\min_{\tau\in\{\tau_{0},\tau_{1},\ldots,\tau_{K}\}}S(\tau;D^{S}).$
This distinction matters. $\tau^{\star}$ is the best admissible script in the full class $\mathcal{T}(C)$. $\hat{\tau}_{K}$ is the best script found by a particular bounded search. The gap between the two reflects the cost of finite search. The audit log records the outcome of every experiment as an ordered sequence $L_{K}=\bigl((k,\tau_{k},s_{k},d_{k})\bigr)_{k=0}^{K}$, where $s_{k}=S(\tau_{k};D^{S})$ when the run succeeds, $s_{k}=+\infty$ when it crashes, and $d_{k}\in\{\textit{keep},\,\textit{discard},\,\textit{crash}\}$ records the outcome status. Algorithm 1 summarizes the protocol.
Algorithm 1 Auditable agent-loop protocol
1:Instruction contract $C$, evaluator $S(\cdot\,;D^{S})$, search sample $D^{S}$, initial script $\tau_{0}$, budget $K$
2:Run evaluator on $\tau_{0}$; record score; set $\tau^{\mathrm{best}}\leftarrow\tau_{0}$
3:for $k=1,\ldots,K$ do
4:  Agent proposes candidate script $\tau_{k}$, informed by log $L_{k-1}$
5:  Run evaluator on $\tau_{k}$; record score (or crash)
6:  If score improves, set $\tau^{\mathrm{best}}\leftarrow\tau_{k}$; otherwise revert to $\tau^{\mathrm{best}}$
7:  Append $(k,\,\tau_{k},\,s_{k},\,\text{outcome})$ to log $L_{k}$
8:end for
The protocol does not impose a model of how the sequence of candidates is generated; that depends on the underlying large language model and the coding environment. Because the agent observes what it has already tried, the search is path dependent and stochastic: independent runs from the same starting point will generally follow different paths and may reach different solutions. The search is adaptive but bounded. The agent searches freely within the editable script, but the evaluator, search sample, and scoring rule are fixed ex ante. This separation between immutable search design and editable implementation is analogous to a pre-analysis plan, except that the audit trail is generated automatically and covers every specification the agent considered. Implementation details, including the instruction contract, are in the Appendix.
Once the search is complete, the researcher may evaluate the selected script $\hat{\tau}_{K}$ on the holdout $D^{H}$. Let $S^{H}(\tau;D^{H})$ denote a scoring function computed on the holdout. The realized holdout score $S^{H}(\hat{\tau}_{K};D^{H})$ provides an external check on whether improvements discovered during adaptive search on $D^{S}$ generalize beyond the search sample. Because $D^{H}$ need not even exist during the search, the holdout evaluation is a researcher action, not part of the agentic loop.
3 Empirical illustration: real-time tuning of peLASSO
3.1 Setup
We illustrate the protocol using the real-time tuning problem in DieboldShin2019’s (DieboldShin2019) peLASSO forecast combination. peLASSO first uses LASSO to select a subset of forecasters from the Survey of Professional Forecasters and then shrinks the surviving weights toward equality using an egalitarian penalty. In the two-step implementation, two regularization parameters must be chosen at each forecast origin using only information available at that time. DieboldShin2019 evaluate the procedure with ex post optimal parameters and do not provide a data-based rule for real-time selection. Instead, they propose other methods such as subset averaging, inspired by the ex post analysis of peLASSO, that perform well on their evaluation sample. The present application takes the tuning problem of peLASSO, which was left unresolved, as the object of agentic search.
The instruction contract asks the agent to find a look-ahead-free lambda selection rule for peLASSO, editing only a single forecasting script (train.R). The immutable evaluator (prepare.R) implements rolling-origin RMSE evaluation under the original DieboldShin2019 design: at each date $t$, only data through t1t-1 is available, and the score is the RMSE over the full evaluation period. We conduct three independent runs, each starting from the simple average as the initial script and an approximate budget of $K=200$ experiments.
We also evaluate the discovered methods on a post-search holdout (2017Q1–2025Q4, 36 quarters) that was withheld from the evaluator during the agentic search. As additional benchmarks, we include two look-ahead-free methods from DieboldShin2019: best average flex ($N_{\max}=6$, their Table 4) and best $(N,W)$ average adaptive (their Table 5), both of which exhaustively search over forecaster subsets at each forecast origin. Details on the holdout data construction are in Appendix E.
3.2 Search-sample and holdout results
Table 1 reports RMSE for the search sample (1999Q3–2016Q4) and the holdout (2017Q1–2025Q4). On the search sample, all three agent runs beat both the simple average and the peLASSO ex post benchmark: Run 1 achieves a relative RMSE of 0.858, Run 3 reaches 0.808, and Run 2 achieves 0.510. The two benchmark methods from the original paper also outperform the simple average, with relative RMSEs of 0.954 (best average flex) and 0.916 (best $(N,W)$ average adaptive).
Table 1: Search-sample and holdout evaluation
Search sample Holdout Holdout excl. COVID
RMSE Relative RMSE Relative RMSE Relative
Simple average $1.504$ $1.000$ $2.979$ $1.000$ $1.120$ $1.000$
peLASSO ex post (original) $1.400$ $0.930$ 3.172 [.880] $1.065$ 1.360 [.885] $1.214$
peLASSO ex post (per window) $1.400$ $0.930$ 2.964 [.191] $0.995$ 1.075 [.136] $0.960$
Best avg flex $1.435$ $0.954$ 2.901 [.240] $0.974$ 1.043 [.347] $0.932$
Best $(N,W)$ avg $1.378$ $0.916$ 2.901 [.127] $0.974$ 1.142 [.602] $1.020$
Run 1 $1.291$ $0.858$ 2.816 [.177] $0.945$ 0.965 [.234] $0.861$
Run 2 $0.767$ $0.510$ 2.417 [.127] $0.811$ 0.827 [.108] $0.739$
Run 3 $1.216$ $0.808$ 3.244 [.852] $1.089$ 1.305 [.755] $1.165$
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Best avg flex and best $(N,W)$ avg are from DieboldShin2019, Tables 4 and 5. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) uses $\lambda$ selected on the search sample for all windows; peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately. Full results including intermediate methods are in Appendix E.
On the holdout, Run 2 remains the strongest method, with a relative RMSE of 0.811 (0.739 excluding the COVID quarters of 2020). Run 1 also holds up, at 0.945 (0.861 excluding COVID). The two benchmark methods from the original paper generalize modestly, each at 0.974. In contrast, Run 3 (1.089) and the peLASSO ex post original (1.065) perform worse than the simple average on the holdout; even the per-window variant that re-optimizes $\lambda$ on each holdout window achieves only 0.995, indicating that peLASSO’s difficulty on the holdout is not merely a lambda selection problem. The holdout therefore shows that agentic search can uncover methods that generalize, but it also reveals that some discovered improvements are sample specific.
3.3 What the log reveals
The instruction contract asks the agent to tune regularization parameters for peLASSO. The experiment logs [^2] Appendix C reproduces an excerpt from the experiment log of Run 3, and Appendix D gives a detailed description of the method discovered by each run., however, reveal that two of the three runs drifted beyond that task: Run 1 toward stability selection and performance weighting, and Run 2 toward a ranking-based method with bias correction. Run 3 stayed closest to the original task, using adaptive LASSO with forward cross validation for lambda selection, though it added an egalitarian elastic net blend. The holdout makes these departures evaluable. Run 2’s bias correction, which exploits serial correlation in combined forecast errors, turns out to generalize and accounts for most of its holdout gain. Run 3’s egalitarian elastic net blend, which drove much of its search-sample improvement, degrades holdout performance. The protocol does not prevent the agent from leaving the intended design space; it ensures that every departure is recorded, so that a holdout or other external evidence can later distinguish structural insight from sample adaptation.
4 Conclusion
Agent loops can widen researcher degrees of freedom, but they need not do so invisibly. This paper adapts a minimal agent loop architecture to an auditable protocol for agentic specification search and adds a post-search holdout evaluation, requiring only a written instruction contract, a fixed evaluator, a single editable script, a finite search budget, and a complete experiment log. In the forecast combination illustration based on DieboldShin2019, three independent runs improve on the search sample, but a genuine holdout shows that only two of the three discovered methods generalize. The contribution of the protocol is not to eliminate adaptive search, but to make it visible, bounded, and evaluable. Because the protocol disciplines the search process rather than the subject matter, it applies wherever automated specification search is used. Appendix F sketches how the same architecture maps to regression specification search, structural VAR identification, and other standard empirical tasks.
Several limitations deserve emphasis. First, a language model may encode background knowledge that overlaps with the holdout period, so residual look-ahead concerns remain. Second, the model’s internal trade-offs between score improvement and contract boundaries are opaque; it is often unclear how those objectives are balanced. Third, the effective search class may recombine existing ideas more readily than it generates novel ones. Fourth, the current search over $\mathcal{T}(C)$ is greedy: propose one candidate, evaluate it, then keep or discard it. More structured strategies, such as the evolutionary population-based search in AlphaEvolve (NovikovEtAl2025) or tree search in AygunEtAl2025, could explore the space more efficiently. Fifth, the protocol is a practical guardrail against data-adaptive search, but it is not a substitute for formal post-selection inference or other corrections for p-hacking.
Appendix A Implementation details
A.1 Four-file architecture
The protocol requires only four files in a shared directory:
$C\equiv\texttt{program.md},\quad\tau\equiv\texttt{train.R},\quad S\equiv\texttt{prepare.R},\quad L_{K}\equiv\texttt{results.tsv}.$
program.md is a plain-text instruction contract that defines the research objective, the admissible modifications, the rules, and the experiment loop of Algorithm 1. train.R is the editable script containing one function the agent rewrites. prepare.R is the immutable evaluator: it loads the search sample, sources train.R, and returns the scalar score; the holdout data is entirely outside the agent’s workspace. results.tsv is a tab-separated audit log in which each row records a commit identifier, the score, the outcome status, and a verbal description of the strategy attempted. In the working directory, the git commit history preserves the actual code of every retained $\tau_{k}$. [^3] The replication archive includes the TSV logs and final code but not the full git object store. Full recoverability from commit history requires access to the original working directory.
The four-file architecture is adapted from Karpathy2026’s (Karpathy2026) autoresearch repository. [^4] See https://github.com/karpathy/autoresearch. The architecture is representative of a broader class of agent loops in which a language model acts as the optimizer within a fixed evaluation harness. The contribution here is not the loop itself but its translation into a protocol for empirical economics: the immutable evaluator fixes what counts as success, the editable script defines the search surface, the instruction contract specifies the researcher’s permissions, and the audit log turns every attempted specification into a reportable object.
A.2 Triggering the agent loop
In the simplest case, the researcher opens a general-purpose coding assistant and prompts it to implement program.md. The agent reads the instruction contract, which describes the experiment loop of Algorithm 1 in plain language, and begins the edit–evaluate–log–retain/revert cycle autonomously.
In our application, we used an outer shell script (run.sh) to restart agent sessions when the model reached its context window limit. Each restart reads results.tsv and the git history to recover state, and the session prompt injects the remaining experiment budget: “You have a budget of $K_{\mathrm{remaining}}$ more experiments. Stop after that.” This is how the budget $K$ is communicated and enforced in practice. The outer script also provides an approximate stop by counting rows in results.tsv and refusing to launch a new session once the count reaches $K$; a session already in progress may overshoot. The full run.sh is available in the replication archive.
To run the protocol, the researcher opens a general-purpose AI coding assistant, such as Claude Code or Codex, and asks it to implement program.md. Because the instruction contract describes the experiment loop in plain language, the agent follows the protocol autonomously. No custom software is needed beyond the four files and a standard coding assistant.
A.3 Isolating the holdout
The post-search holdout $D^{H}$ must remain entirely outside the agent’s workspace during the search. In our application, the holdout period (2017Q1–2025Q4) was constructed after the agentic search was complete, so no holdout data existed in the working directory while the agent was running. More generally, researchers should ensure that holdout data is not accessible to the agent during the search. Because current coding agents can access the internet and execute arbitrary code, an agent could in principle download or construct holdout period data on its own. Disabling internet access during the search, restricting the agent’s file system permissions, or simply withholding the holdout data from the working directory are practical safeguards. Without such boundaries, the transparency of the holdout evaluation cannot be guaranteed.
Appendix B Instruction contract
The following is the verbatim program.md used in the empirical application of Section 3.
# autoresearch  PE Lasso lambda selection
## Goal
Find the best look-ahead-free, data-based lambda selection algorithm
for PE Lasso forecast combination.
PE Lasso forecast combination we mean partially egalitary lasso method
in the paper /original_paper/DiebodShinEgalitarianLasso.md
(both one-step conceptualization and two-step implementation).
For both cases, we have two regularization parameters.
IMPORTANT: Focus on two-step implementation of PE Lasso. Do not try
to combine with other methods. I want to focus on finding a way to
choose these regularization parameters for PE Lasso so that it becomes
tuning free method. We could try bunch of different CV methods. Or,
other data adaptive method automatically selects these tuning
parameters. We understand it is difficult problem finding two
regularization parameters from two different stages. But, this is
exactly why we want to run this experiments.
The simple average benchmark has RMSE ~1.50. The ex-post oracle
(which cheats by using future data) achieves ~1.40. Your job is to
close that gap.
You decide what strategies to explore. The reference implementation
is in ‘original_program/R/‘  read it for ideas.
## Setup
1. Agree on a run tag with the user: propose a tag based on today’s
   date (e.g. ‘mar11‘). The branch ‘autoresearch/<tag>‘ must not
   already exist.
2. Create the branch: ‘git checkout -b autoresearch/<tag>‘
3. Read the in-scope files:
   - ‘program.md‘  this file (your instructions)
   - ‘prepare.R‘  fixed evaluation harness. Do not modify.
   - ‘train.R‘  the file you modify. Defines ‘select_lambda(info)‘.
4. Check session history: Read ‘results.tsv‘ and ‘git log‘ to
   understand what has been tried before. This is your memory across
   sessions.
5. Initialize results.tsv if it doesn’t exist: create it with just
   the header row.
6. Confirm and go.
## The rules
- One file: You only edit ‘train.R‘. It must define
  ‘select_lambda(info)‘ returning ‘list(forecast, method)‘.
- No lookahead: At time t, only data from 1:(t-1) is available.
  The ‘info‘ object enforces this.
- No new packages: Only glmnet, forecast, and base R.
- Runtime: Each full run (66 evaluation periods) should complete
  in under 30 minutes.
- Deterministic: Use a fixed seed if randomness is needed.
## What’s in ‘info‘
  info\$X_train      # rolling window training forecasts [max(1,t-w):(t-1), 23]
  info\$y_train      # corresponding actuals
  info\$X_history    # full history [1:(t-1), 23]
  info\$y_history    # full history actuals
  info\$x_new        # current period forecasts [23]
  info\$lambda_grid  # 200-point lambda grid, exp(15) to exp(-15)
  info\$lambda_grid2 # same grid (for two-step methods)
  info\$t            # current time index
  info\$w            # window size (20)
All functions from ‘prepare.R‘ are in scope: ‘pelasso_lasso_avg()‘,
‘pelasso_lasso_eridge()‘, ‘pelasso_lasso_elasso()‘, ‘fit_lasso()‘,
‘fit_ridge()‘, ‘fit_elasso()‘, ‘fit_eridge()‘,
‘cv_loo_bandwidth()‘, ‘cv_loo_bandwidth_2step()‘, etc.
## Benchmarks
| Method                               | RMSE  | Notes                      |
| Simple average                       | ~1.50 | Baseline (no lambda)       |
| Ex-post optimal peLASSO(LASSO, Avg)  | ~1.40 | Oracle  uses future data  |
| Good data-based method               | 1.42-1.48 | Realistic target range |
## Output format
The script prints:
  method:         <name>
  rmse:           <value>
  benchmark_rmse: <simple average RMSE>
  relative_rmse:  <ratio>
Extract the key metric: ‘grep "^rmse:" run.log‘
## Logging results
Log every experiment to ‘results.tsv‘ (tab-separated).
Do NOT commit this file.
  commit   rmse    status   description
  a1b2c3d  1.5012  keep     baseline simple average
  b2c3d4e  1.4723  keep     LOO CV for pelasso_lasso_avg
  c3d4e5f  1.4891  discard  BIC-based lambda selection
  d4e5f6g  0.0000  crash    2D CV OOM on full grid
## Session recovery
You may be starting a fresh session after a previous one was
interrupted (context limit, crash, etc.). On every session start:
1. Read ‘results.tsv‘ and ‘git log --oneline‘ to see what has
   been tried.
2. Reconcile state: Compare the last ‘results.tsv‘ entry to
   the latest git commit.
   - If the latest commit is NOT in ‘results.tsv‘, it was never
     evaluated. Run it now and log the result before moving on.
   - If ‘train.R‘ has uncommitted changes, they are leftover from
     an interrupted edit. Review and either commit or discard.
3. Identify the current best RMSE from ‘results.tsv‘ (the lowest
   ‘rmse‘ with status ‘keep‘).
4. Continue experimenting from there.
## The experiment loop
LOOP (until experiment budget is exhausted or interrupted):
1.  Read the current state: ‘train.R‘, ‘results.tsv‘, git log
2.  Edit ‘train.R‘ with a new lambda selection strategy
3.  ‘git commit -m "descriptive message"‘
4.  Run: ‘Rscript prepare.R > run.log 2>&1‘
5.  Extract: ‘grep "^rmse:" run.log‘
6.  If empty (crash): ‘tail -n 50 run.log‘ to diagnose
7.  Log to ‘results.tsv‘
8.  If RMSE improved -> keep the commit
9.  If RMSE worse -> ‘git reset --hard HEAD~1‘
10. Repeat
Logging is step 7, BEFORE the keep/revert decision in steps 8-9.
This ensures every experiment is recorded even if the session is
interrupted right after.
Timeout: If a run exceeds 30 minutes, kill it and treat as crash.
Crashes: Fix trivial bugs (typos, off-by-one). If the idea is
fundamentally broken, log crash and move on.
NEVER STOP: Do not pause to ask if you should continue. Run
autonomously until your experiment budget runs out or you are
manually interrupted. If you run out of ideas, re-read
‘original_program/R/‘ for new angles, try combining previous
near-misses, or try more radical approaches.
Appendix C Experiment log excerpt
Table 2 presents a selected excerpt from the 201-experiment log of Run 3. The trajectory illustrates a systematic exploration pattern. The agent begins with standard cross-validation for peLASSO (experiments 1–4), quickly moves to adaptive LASSO with ridge-based penalty weights and forward cross-validation (experiments 25–45), discovers that multi-horizon forward CV improves regularization parameter selection (experiment 86), introduces model averaging over near-optimal lambda values (experiment 113), and finally arrives at an ensemble that blends peLASSO with an egalitarian elastic net component (experiments 138–201). Each phase builds on the structural lessons of the previous one.
Table 2: Experiment log excerpt: Run 3
Exp. Status RMSE Description
1 keep $1.5043$ baseline simple average
2 keep $1.4692$ pelasso(lasso,avg) LOO CV B=0
3 crash pelasso(lasso,avg) LOO CV B=1 – zero variance error
4 discard $1.9556$ pelasso(lasso,eridge) 2D LOO CV coarse 30pt grid
25 keep $1.4455$ pe(adalasso,avg) fwdCV exp 0.95 ridge penalty weights
45 keep $1.3839$ pe(adalasso,avg) rolling window ridge penalties decay=0.6
86 keep $1.3539$ pe(adalasso,avg) h=2 step ahead forward CV LOO ridge
113 keep $1.3202$ pe(adalasso,avg) h=2 model averaging top 2pct lambdas
138 keep $1.2533$ pe(adalasso,blend) 80pct PE + 20pct eridge ensemble
152 keep $1.2253$ pe(adalasso,blend) 70/30 PE + eEN cv.glmnet LOO
201 keep $1.2159$ pe(adalasso,blend) model avg 1pct h1(0.05)+h2(0.95)
  • Notes: Selected rows from the 201-experiment log of Run 3. Descriptions are the agent’s verbatim one-line summaries, truncated for display. Experiment 1 is the initial simple average baseline. RMSE for crash entries is undefined; the implementation records them as zero in results.tsv, but they are excluded from all comparisons (formally, $s_{k}=+\infty$). Full logs are in the replication archive.
Appendix D Discovered methods
Each agent run produced a final method and several intermediate variants. The holdout evaluation in Section E reports results for the final method and two representative intermediates per run, selected from distinct algorithmic phases. We describe each method here.
Except where noted, all methods operate within the same rolling window framework. At each forecast origin $t$, let $\mathbf{X}_{t}\in\mathbb{R}^{n_{t}\times K}$ and $\mathbf{y}_{t}\in\mathbb{R}^{n_{t}}$ denote the training window of forecaster predictions and GDP realizations, where $n_{t}=\min(t-1,W)$, $W=20$, and $K=23$. Let $\mathbf{x}_{t}=(x_{1,t},\ldots,x_{K,t})^{\prime}$ denote the new forecaster predictions for period $t$. Each method produces a combined forecast $\hat{y}_{t}$; most methods use only $(\mathbf{X}_{t},\mathbf{y}_{t},\mathbf{x}_{t})$, with exceptions noted below.
D.1 Run 1: stability selection with performance weighting
Run 1 (final).
Stability selection with performance weighted aggregation.
  1. 1.
    Fit elastic net ($\alpha=0.65$) with 5 fold CV on the training window. Define $\hat{\lambda}_{1.5\mathrm{se}}$ as the largest $\lambda$ such that the CV error is within $1.5$ standard errors of the minimum.
  2. 2.
    Stability selection. Let $\mathcal{W}=\{j:|\log\lambda_{j}-\log\hat{\lambda}_{1.5\mathrm{se}}|<\delta\}$ with $\delta=0.25$. The selection frequency of forecaster $k$ is
    $\pi_{k}=|\mathcal{W}|^{-1}\sum_{j\in\mathcal{W}}\mathbf{1}\{\hat{\beta}_{k}(\lambda_{j})\neq 0\}.$ (1)
    The active set is $\mathcal{A}=\{k:\pi_{k}>0.4\}$.
  3. 3.
    Composite weights. Using the $m=2$ most recent training observations, compute individual $\mathrm{RMSE}_{k}$ for each $k\in\mathcal{A}$ and form
    $w_{k}=\pi_{k}^{2}\cdot\mathrm{RMSE}_{k}^{-p},\quad p=14.$ (2)
  4. 4.
    Dominance check. If $w_{(1)}>5\cdot w_{(2)}$ (where $w_{(j)}$ denotes the $j$th largest weight), use the single best forecaster.
  5. 5.
    Otherwise, the combined forecast is the weighted $q$th quantile ($q=0.44$) of $\{x_{k,t}\}_{k\in\mathcal{A}}$ with normalized weights $w_{k}/\sum_{k}w_{k}$, computed by linear interpolation of the weighted empirical CDF.
Run 1a (Phase 2: elastic net with median).
PE selection using elastic net ($\alpha=0.65$) with 5 fold CV and $\hat{\lambda}_{1\mathrm{se}}$. The combined forecast is the median of the selected forecasters:
$\hat{y}_{t}=\mathrm{median}\bigl(\{x_{k,t}\}_{k\in\mathcal{A}(\hat{\lambda}_{1\mathrm{se}})}\bigr).$ (3)
No stability selection, no performance weighting.
Run 1b (Phase 1: peLASSO with CV).
Standard peLASSO(LASSO, Avg) with $\hat{\lambda}_{1\mathrm{se}}$ from cv.glmnet:
$\hat{y}_{t}=|\mathcal{A}(\hat{\lambda}_{1\mathrm{se}})|^{-1}\sum_{k\in\mathcal{A}(\hat{\lambda}_{1\mathrm{se}})}x_{k,t}.$ (4)
This is the closest variant to the original DieboldShin2019 method, but with data driven $\lambda$ selection rather than ex post optimization.
D.2 Run 2: ranking, adaptive windows, and bias correction
Run 2 (final).
Temporally weighted ranking with adaptive window selection and bias correction.
For a given training window $(\mathbf{X},\mathbf{y})$ of size $n$:
  1. 1.
    Temporal weights. Define $\omega_{s}=\exp\bigl(-\alpha(n-s)/(n-1)\bigr)/\sum_{j=1}^{n}\exp\bigl(-\alpha(n-j)/(n-1)\bigr)$ with $\alpha=6$. More recent observations receive higher weight.
  2. 2.
    Weighted MAE ranking. $\mathrm{MAE}_{k}=n^{-1}\sum_{s=1}^{n}n\omega_{s}|y_{s}-x_{k,s}|$. Rank forecasters by ascending MAE.
  3. 3.
    LOO CV for subset size. For each candidate $N\in\{3,\ldots,\min(K,18)\}$ and each left out observation $i$:
    • Rerank by RMSE on the remaining n1n-1 observations.
    • Select the top $N$ forecasters, weight them by $\tilde{w}_{j}\propto\mathrm{MAE}_{j}^{-3}$ (on leave out data).
    • Compute the weighted average LOO forecast $\hat{y}_{i}^{\mathrm{LOO}}=\sum_{j}\tilde{w}_{j}x_{j,i}$ and the median based LOO forecast $\hat{y}_{i}^{\mathrm{pct}}=Q_{0.50}(\{x_{j,i}\}_{j\in\text{top }N})$.
    Select $N^{*}=\arg\min_{N}\sqrt{n^{-1}\sum_{i}(\hat{y}_{i}^{\mathrm{LOO}}-y_{i})^{2}}$.
  4. 4.
    Adaptive window. Evaluate steps 1–3 at the full rolling window and at each sub-window $w^{\prime}\in\{4,5,\ldots,19\}$, keeping whichever $(N^{*},w^{\prime})$ combination yields the lowest LOO RMSE.
  5. 5.
    Median aggregation. The base forecast is $\tilde{y}_{t}=Q_{0.50}(\{x_{k,t}\}_{k\in\text{top }N^{*}})$.
  6. 6.
    Bias correction. The final forecast adjusts for serial correlation in forecast errors:
    $\hat{y}_{t}=\tilde{y}_{t}-\gamma\cdot\hat{e}_{n}^{\mathrm{LOO,pct}},\quad\gamma=0.80,$ (5)
    where $\hat{e}_{n}^{\mathrm{LOO,pct}}=\hat{y}_{n}^{\mathrm{pct}}-y_{n}$ is the percentile based LOO error from the most recent training observation.
Run 2a (no bias correction).
Identical to Run 2 (steps 1–5) but omits the bias correction (5):
$\hat{y}_{t}=\tilde{y}_{t}.$ (6)
This isolates the contribution of the forecaster ranking and adaptive window mechanism.
Run 2b (weighted ranking, no adaptive window).
Steps 1–2 as in Run 2 with a fixed window. Step 3 uses a penalized LOO criterion, $\mathrm{RMSE}_{\mathrm{LOO}}(N)+0.01\cdot\log(N+1)$, to select $N^{*}$. The combined forecast is a weighted average (not median) of the top $N^{*}$ forecasters:
$\hat{y}_{t}=\sum_{j=1}^{N^{*}}\tilde{w}_{j}\,x_{\sigma(j),t},\quad\tilde{w}_{j}\propto\mathrm{MAE}_{\sigma(j)}^{-3},$ (7)
where $\sigma$ denotes the MAE ranking. No adaptive window, no bias correction.
D.3 Run 3: adaptive LASSO with forward CV and ensemble
Run 3 (final).
Adaptive LASSO with multi horizon forward CV, model averaging, and egalitarian elastic net ensemble.
  1. 1.
    Pilot estimator. Fit elastic net ($\alpha_{\mathrm{pilot}}=0.1$) with LOO CV on the rolling window to obtain $\hat{\boldsymbol{\beta}}^{\mathrm{ridge}}$.
  2. 2.
    Adaptive LASSO penalties. Define forecaster specific penalty factors
    $\tilde{w}_{k}=\frac{(|\hat{\beta}_{k}^{\mathrm{ridge}}|+\epsilon)^{-1}}{\overline{(|\hat{\beta}^{\mathrm{ridge}}|+\epsilon)^{-1}}},\quad\epsilon=5\times 10^{-3},$ (8)
    where $\overline{\cdot}$ denotes the cross sectional mean. Fit LASSO with penalty $\lambda\cdot\tilde{w}_{k}|\beta_{k}|$ on the full history.
  3. 3.
    Multi horizon forward CV. For each fold $f$ with training end $t_{f}$, compute forecast errors at horizons $h=1$ and $h=2$. Two candidate aggregation rules are evaluated: simple average and blended average ($0.7\times\text{mean}+0.3\times\text{ridge weighted mean}$). The CV criterion is
    $\mathrm{CV}(\lambda)=\sum_{f}w_{f}\left(0.05\cdot e_{f,h=1}^{2}(\lambda)+0.95\cdot e_{f,h=2}^{2}(\lambda)\right),$ (9)
    with fold weights $w_{f}\propto\gamma^{F-f}$ and $\gamma=0.75$.
  4. 4.
    Model averaging. Average forecasts over all $\lambda$ within 1% of the best CV score, weighted by inverse RMSE.
  5. 5.
    Egalitarian elastic net. Fit elastic net ($\alpha=0.5$) via cv.glmnet with LOO CV on the rolling window, selecting $\hat{\lambda}_{\min}$, on the egalitarian transformation $\tilde{y}_{s}=y_{s}-K^{-1}\sum_{k}x_{k,s}$, obtaining $\tilde{\boldsymbol{\beta}}^{\mathrm{eEN}}$, and compute $\hat{y}_{t}^{\mathrm{eEN}}=\sum_{k}(\tilde{\beta}_{k}^{\mathrm{eEN}}+K^{-1})\,x_{k,t}$.
  6. 6.
    Final blend.
    $\hat{y}_{t}=0.70\cdot\hat{y}_{t}^{\mathrm{PE}}+0.30\cdot\hat{y}_{t}^{\mathrm{eEN}}.$ (10)
Run 3a (Phase 4: single horizon, no blend).
Steps 1–2 as in Run 3. Forward CV with $h=1$ only and $\gamma=0.75$. No model averaging. Simple average of selected forecasters. No egalitarian elastic net blend:
$\hat{y}_{t}=|\mathcal{A}(\hat{\lambda})|^{-1}\sum_{k\in\mathcal{A}(\hat{\lambda})}x_{k,t}.$ (11)
Run 3b (Phase 5: multi horizon, no blend).
Steps 1–4 as in Run 3, except $\gamma=0.80$ and $\epsilon=10^{-3}$. Model averaging is applied. No egalitarian elastic net blend: $\hat{y}_{t}=\hat{y}_{t}^{\mathrm{PE}}$.
Appendix E Post-search holdout evaluation
This section reports the full holdout evaluation summarized in Section 3 of the main text. The holdout period (2017Q1–2025Q4, 36 quarters) was withheld from the evaluator during the agentic search.
E.1 Data and processing
The extended dataset combines ECB Survey of Professional Forecasters (SPF) individual microdata with Eurostat GDP realizations (table namq_10_gdp, chain linked volumes, seasonally and calendar adjusted). The original sample (1999Q3–2016Q4, 70 observations) is copied verbatim from the original H1_gdp.csv of DieboldShin2019 to guarantee exact RMSE reproduction. The new sample (2017Q1–2025Q4, 36 quarters) is constructed from the SPF extraction. Missing forecaster responses in the new sample are imputed by linear interpolation for interior gaps and by cross-sectional mean for trailing gaps and inactive forecasters. COVID quarters are defined as 2020Q1–2020Q4.
E.2 Experiment counts
Table 3 reports the experiment counts for the three agent runs. Each run uses Claude Code with Opus 4.6 at default thinking effort, starting from a fresh agent with the same initial script (simple average) and an approximate budget of $K=200$ experiments. The evaluator calls the forecasting function 66 times ($t=5,\ldots,70$) but scores RMSE over 65 periods ($t=6,\ldots,70$); the first call serves as a warm-up. The budget is enforced as a soft constraint; the realized counts are 232, 228, and 201 for Runs 1, 2, and 3, respectively. If a candidate script exceeds 30 minutes of runtime, the experiment is terminated and logged as a crash.
Table 3: Experiment counts
Experiments Keeps Discards Crashes
Run 1 232 42 188 2
Run 2 228 61 164 3
Run 3 201 46 152 3
  • Notes: “Keep” means the modification improved on the current best RMSE; “discard” means it did not; “crash” means the candidate script exceeded the 30-minute timeout or produced an error. The budget is enforced as a soft constraint; realized counts vary.
E.3 Full holdout results
Table 4 reports RMSE for all methods, including intermediate variants that isolate individual components. For each run, we include two representative intermediate methods selected from distinct algorithmic phases. The intermediate methods are informative because they isolate the contribution of individual components such as stability selection, bias correction, or ensemble blending.
Table 4: Full holdout evaluation of forecast combination methods
Search sample Holdout Holdout excl. COVID
RMSE Relative RMSE Relative RMSE Relative
Panel A: Benchmarks and original paper methods
Simple average $1.504$ $1.000$ $2.979$ $1.000$ $1.120$ $1.000$
peLASSO ex post (original) $1.400$ $0.930$ 3.172 [.880] $1.065$ 1.360 [.885] $1.214$
peLASSO ex post (per window) $1.400$ $0.930$ 2.964 [.191] $0.995$ 1.075 [.136] $0.960$
Best avg flex $1.435$ $0.954$ 2.901 [.240] $0.974$ 1.043 [.347] $0.932$
Best $(N,W)$ avg $1.378$ $0.916$ 2.901 [.127] $0.974$ 1.142 [.602] $1.020$
Best individual $1.403$ $0.933$ 2.994 [.716] $1.005$ 1.164 [.714] $1.040$
Panel B: Run 1 family
Run 1 (final) $1.291$ $0.858$ 2.816 [.177] $0.945$ 0.965 [.234] $0.861$
Run 1a (ENet, median) $1.390$ $0.924$ 2.991 [.742] $1.004$ 1.158 [.794] $1.034$
Run 1b (peLASSO, CV) $1.451$ $0.964$ 3.059 [.938] $1.027$ 1.270 [.929] $1.134$
Panel C: Run 2 family
Run 2 (final) $0.767$ $0.510$ 2.417 [.127] $0.811$ 0.827 [.108] $0.739$
Run 2a (no bias corr.) $1.308$ $0.869$ 2.949 [.087] $0.990$ 1.093 [.290] $0.976$
Run 2b (weighted avg) $1.351$ $0.898$ 2.918 [.113] $0.980$ 1.120 [.499] $1.000$
Panel D: Run 3 family
Run 3 (final) $1.216$ $0.808$ 3.244 [.852] $1.089$ 1.305 [.755] $1.165$
Run 3a (h=1h{=}1, no blend) $1.309$ $0.870$ 2.994 [.648] $1.005$ 1.238 [.818] $1.106$
Run 3b (h=2h{=}2, no blend) $1.307$ $0.869$ 3.037 [.865] $1.020$ 1.276 [.855] $1.140$
  • Notes: RMSE of forecast errors for euro area real GDP growth (year on year). Search sample: 1999Q3–2016Q4 (65 evaluation periods). Holdout: 2017Q1–2025Q4 (36 quarters). Holdout excl. COVID: holdout dropping 2020Q1–Q4 (32 quarters). Relative RMSE is computed against the simple average on each respective sample. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test that the method has superior predictive accuracy relative to the simple average, using the EWC fixed-$b$ approximation (ShinSchor2026); these have intrinsically low power given the small holdout sample (36 quarters, 32 excluding COVID) and serial correlation in forecast errors, and should be read as descriptive. peLASSO ex post (original) and best individual use parameters selected on the search sample and are evaluated without re-optimization on the holdout. peLASSO ex post (per window) re-optimizes $\lambda$ on each evaluation window separately.
E.4 Role of individual components
The intermediate methods isolate what drives generalization.
Stability selection (Run 1 vs. 1a, 1b).
The full Run 1 (holdout relative RMSE 0.945) substantially outperforms both alternatives: Run 1a (1.004), which uses elastic net selection without stability, and Run 1b (1.027), which uses standard peLASSO with CV. The stability selection mechanism and the composite performance weighting are jointly responsible for the holdout gain. Without them, elastic net PE selection is essentially no better than the simple average.
Bias correction (Run 2 vs. 2a, 2b).
Removing the bias correction from Run 2 degrades performance from 0.811 to 0.990 (Run 2a) or 0.980 (Run 2b). The bias correction contributes most of Run 2’s advantage. However, Run 2a and 2b are still competitive with the simple average, indicating that the temporally weighted ranking and adaptive window mechanism also add modest value. The bias correction exploits a structural feature—serial correlation in combined forecast errors—rather than overfitting to the search sample.
Egalitarian elastic net blend (Run 3 vs. 3a, 3b).
Run 3’s 70/30 blend with the egalitarian elastic net (holdout relative RMSE 1.089) performs worse than the simpler PE adaptive LASSO without the blend (3a at 1.005, 3b at 1.020). The ensemble component that improved search-sample fit introduced additional estimation noise that does not pay off on the holdout.
E.5 Independent dataset robustness
To verify that these results are not an artifact of using the original H1_gdp.csv verbatim for the search-sample period, we also constructed a fully independent dataset where all 106 rows are built from downloaded SPF and Eurostat data, with no verbatim copy from the original file. The independent dataset uses the current (2026) GDP vintage rather than the 2018Q1 vintage in the original study; over the original sample the mean absolute revision is 0.079 pp and the maximum is 0.20 pp.
Table 5 repeats the exercise on this independent dataset. The key rankings are preserved: Run 2 achieves a relative RMSE of 0.808 on the holdout, Run 1 achieves 0.946, and the original paper methods achieve 0.974 and 0.968. The only notable difference is Run 3, which improves from 1.089 to 0.962, suggesting its search-sample performance was sensitive to the specific GDP vintage. The independent dataset confirms that the main ranking—Run 2 strongest, Run 1 second—is preserved across vintages, while Run 3 is sensitive to the GDP vintage used.
Table 5: Full holdout evaluation: independent dataset
Search sample Holdout Holdout excl. COVID
RMSE Relative RMSE Relative RMSE Relative
Panel A: Benchmarks and original paper methods
Simple average $1.510$ $1.000$ $2.979$ $1.000$ $1.120$ $1.000$
peLASSO ex post (original) $1.417$ $0.938$ 3.041 [.763] $1.021$ 1.350 [.881] $1.206$
peLASSO ex post (per window) $1.417$ $0.938$ 2.976 [.249] $0.999$ 1.119 [.188] $0.999$
Best avg flex $1.500$ $0.994$ 2.901 [.239] $0.974$ 1.043 [.346] $0.931$
Best $(N,W)$ avg $1.298$ $0.860$ 2.884 [.100] $0.968$ 1.094 [.340] $0.977$
Best individual $1.407$ $0.932$ 2.909 [.245] $0.977$ 1.109 [.463] $0.990$
Panel B: Run 1 family
Run 1 (final) $1.295$ $0.858$ 2.817 [.177] $0.946$ 0.965 [.233] $0.861$
Run 1a (ENet, median) $1.479$ $0.980$ 2.991 [.730] $1.004$ 1.157 [.779] $1.033$
Run 1b (peLASSO, CV) $1.455$ $0.964$ 3.051 [.952] $1.024$ 1.250 [.931] $1.116$
Panel C: Run 2 family
Run 2 (final) $0.834$ $0.552$ 2.408 [.127] $0.808$ 0.826 [.107] $0.737$
Run 2a (no bias corr.) $1.304$ $0.864$ 2.947 [.090] $0.989$ 1.096 [.298] $0.978$
Run 2b (weighted avg) $1.343$ $0.890$ 2.898 [.100] $0.973$ 1.065 [.054] $0.950$
Panel D: Run 3 family
Run 3 (final) $1.305$ $0.864$ 2.865 [.309] $0.962$ 1.349 [.728] $1.205$
Run 3a (h=1h{=}1, no blend) $1.490$ $0.987$ 2.960 [.400] $0.994$ 1.270 [.802] $1.133$
Run 3b (h=2h{=}2, no blend) $1.502$ $0.995$ 2.963 [.414] $0.995$ 1.276 [.812] $1.139$
  • Notes: Same as Table 4 but using the independently constructed dataset (all rows from downloaded SPF and Eurostat data, no verbatim copy from the original H1_gdp.csv). Search-sample RMSEs differ from Table 4 due to GDP vintage differences and independent imputation. Bracketed values in the holdout columns are one sided $p$-values from the Diebold–Mariano test using the EWC fixed-$b$ approximation (ShinSchor2026); see Table 4 notes for details and caveats. peLASSO ex post (original) and best individual use parameters selected on the search sample; see Table 4 notes for details.
Appendix F Illustration: forecast combination and beyond
In the empirical application of DieboldShin2019 studied in Section 3, program.md asks the agent to find a look-ahead-free, data-based lambda selection rule for peLASSO forecast combination. train.R defines a function that takes admissible historical data and returns a forecast. prepare.R implements a rolling-origin evaluation: at each date $t$, it passes train.R only information available through (t1)(t-1), produces a forecast for $t$, and computes RMSE over the evaluation period. results.tsv records one line per candidate rule. The empirical task is narrow and inspectable: construct a function that maps admissible historical information into one forecast, and judge every attempted rule under the same rolling-origin RMSE criterion.
The same four-file architecture applies whenever the researcher can separate an immutable evaluation design from an editable implementation rule. The remainder of this section provides three stylized illustrations that map the protocol into familiar empirical settings: regression specification search, structural VARs, and inflation forecasting for monetary policy.
These illustrations are intentionally schematic. In forecasting problems, the evaluator can often be a straightforward out-of-sample loss. In inference-heavy settings, by contrast, any scalar score should be understood as a diagnostic screen within a broader research design, not as a complete validation of identification or inference.
Example 1: Specification search in a regression model.
Consider a researcher studying a descriptive wage regression using cross-sectional data. The estimating equation is
$\log w_{i}\;=\;\beta\cdot\text{educ}_{i}+g(z_{i})+e_{i},$
where $w_{i}$ is earnings, $\text{educ}_{i}$ is years of schooling, $\beta$ is a coefficient of interest, and $z_{i}$ is a vector of controls including experience, experience squared, demographic indicators, occupation, industry, and geographic region. The researcher wants to compare alternative specifications for the nuisance function $g(\cdot)$ while handling inference on $\beta$ separately.
File mapping.
  • program.md ($C$): “Compare candidate specifications of $g(z_{i})$ using a held-out evaluation sample. Only edit train.R. The evaluation sample partition is fixed in prepare.R. Do not modify the sample or the outcome variable.”
  • train.R ($\tau$): Defines a function that receives the training sample and returns fitted values for the nuisance component, potentially along with fitted values for $\text{educ}_{i}$ if a partialling-out design is used. The agent may explore polynomial terms, splines, interactions, LASSO-penalized specifications, or other functional forms for $g$.
  • prepare.R ($S$): Splits the data into a training sample $\mathcal{I}_{\text{train}}$ and an evaluation sample $\mathcal{I}_{\text{eval}}$, fixed ex ante. On $\mathcal{I}_{\text{train}}$, it sources train.R to obtain the nuisance fit, and if desired partials out both outcome and treatment before estimating $\hat{\beta}$. On $\mathcal{I}_{\text{eval}}$, it computes a holdout prediction error such as
    $S(\tau;D)=\frac{1}{|\mathcal{I}_{\text{eval}}|}\sum_{i\in\mathcal{I}_{\text{eval}}}\bigl(\log w_{i}-\hat{\beta}\cdot\text{educ}_{i}-\hat{g}(z_{i})\bigr)^{2}.$
  • results.tsv ($L_{K}$): Records each specification attempted, its holdout MSE, and the implied $\hat{\beta}$.
If the scoring rule were the in-sample $p$-value for $\hat{\beta}$, the loop would amount to automated $p$-hacking. The discipline comes from fixing the evaluator and the sample split before the search begins and from disclosing the full search path. Causal claims about returns to schooling would still require separate identification assumptions.
Example 2: Structural VAR identification.
Consider a three-variable VAR for macroeconomic analysis of monetary policy. Let $y_{t}=(\Delta\text{GDP}_{t},\;\pi_{t},\;i_{t})^{\prime}$, where $\Delta\text{GDP}_{t}$ is real GDP growth, $\pi_{t}$ is inflation, and $i_{t}$ is the federal funds rate. The reduced-form VAR is
$y_{t}=\Phi_{1}y_{t-1}+\cdots+\Phi_{p}y_{t-p}+u_{t},$
where $u_{t}$ is the reduced-form innovation with $\text{Var}(u_{t})=\Sigma_{u}$. The structural representation is $u_{t}=A_{0}^{-1}\varepsilon_{t}$, where $\varepsilon_{t}$ is the vector of structural shocks and $A_{0}$ encodes the contemporaneous causal structure. The goal here is more modest: compare candidate ways of constructing a monetary policy shock series and record how each one lines up with an external narrative diagnostic.
File mapping.
  • program.md ($C$): “Compare candidate constructions of a monetary policy shock series using a fixed external diagnostic on a held-out evaluation sample. Only edit train.R. Do not modify the VAR specification or the evaluation sample.”
  • train.R ($\tau$): Defines a function that receives the estimated reduced-form parameters $(\hat{\Phi}_{1},\ldots,\hat{\Phi}_{p},\hat{\Sigma}_{u})$, together with any auxiliary diagnostic series supplied by prepare.R, and returns a candidate monetary policy shock series $\hat{\varepsilon}_{t}^{mp}$ and the associated GDP impulse response at horizon $h$. The agent may explore recursive orderings, sign restrictions, or other mappings from reduced-form innovations into candidate shock measures.
  • prepare.R ($S$): Estimates the reduced-form VAR on a fixed training sample, calls train.R to obtain the candidate shock, and computes a diagnostic score on a held-out evaluation period:
    $S(\tau;D)=-\left|\text{corr}\!\left(\hat{\varepsilon}_{t}^{mp},\;z_{t}^{RR}\right)\right|,$
    where $z_{t}^{RR}$ is the Romer and Romer narrative monetary policy shock measure and the evaluation period is fixed ex ante. The negative sign ensures that higher correlation corresponds to a lower (better) score. This score is best read as an external diagnostic rather than a standalone proof of identification.
  • results.tsv ($L_{K}$): Records the candidate construction, the diagnostic score, and the implied peak response of GDP to a monetary policy shock.
The point is not that the agent “finds the largest effect” or settles identification, but that any search over candidate shock constructions is visible, bounded, and judged against a criterion fixed ex ante.
Example 3: Inflation forecasting for monetary policy.
Central banks routinely produce inflation forecasts to guide monetary policy decisions. The workhorse framework is the Phillips curve:
$\pi_{t+h}=\alpha+\beta\cdot\text{slack}_{t}+\gamma\cdot\pi_{t}^{e}+\delta^{\prime}z_{t}+e_{t+h},$
where $\pi_{t+h}$ is $h$-quarter-ahead CPI inflation, $\text{slack}_{t}$ is a measure of economic slack, $\pi_{t}^{e}$ is an inflation expectations proxy, and $z_{t}$ is a vector of additional conditioning variables (commodity prices, exchange rates, financial conditions). The practical difficulty is that the specification involves many choices: which slack measure (unemployment gap, output gap, capacity utilization), which expectations proxy (survey expectations, breakeven inflation, adaptive expectations), which additional predictors, which estimation method (OLS, ridge, random forest), and which estimation window.
File mapping.
  • program.md ($C$): “Compare candidate real-time Phillips curve specifications for $h=4$ quarter-ahead CPI inflation forecasting. No lookahead: at forecast origin $t$, only real-time vintage data through $t$ is available. Only edit train.R. Budget: $K=200$ experiments.”
  • train.R ($\tau$): Defines a function that receives the information set $\mathcal{F}_{t}=\{y_{s},x_{s}\}_{s=1}^{t}$ and returns a point forecast $\hat{\pi}_{t+h|t}$. The agent may explore different slack measures, expectations proxies, transformations, variable selection methods, penalized estimation, or nonlinear specifications.
  • prepare.R ($S$): Implements pseudo-out-of-sample evaluation. For each forecast origin $t\in\{t_{0},t_{0}+1,\ldots,T-h\}$, passes the real-time information set to train.R and collects the forecast. The score is
    $S(\tau;D)=\text{RMSE}=\left(\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\bigl(\pi_{t+h}-\hat{\pi}_{t+h|t}\bigr)^{2}\right)^{1/2}.$
  • results.tsv ($L_{K}$): Records each Phillips curve specification, its RMSE, and a description of the method (e.g., “ridge Phillips curve, unemployment gap, SPF expectations, 40-quarter rolling window”).
This is a natural fit for the protocol because the evaluator is a genuine pseudo-out-of-sample forecast loss. For central bank staff who routinely compare Phillips curve specifications, the protocol would make the search path more visible and easier to disclose.
These three illustrations suggest how the four-file protocol can be adapted across settings. The architecture applies whenever the researcher can separate an immutable evaluation design from an editable implementation rule. Forecasting is the cleanest instance because the score is naturally out-of-sample. In more inference-heavy settings, the protocol is better viewed as a way to make adaptive search more visible and easier to disclose, not as a substitute for the underlying econometric argument.
A post-search holdout can also help reduce concerns about p-hacking, but only in a limited sense. Its value is that it separates discovery from evaluation: the agent may search adaptively over many specifications on the search sample, but the selected output is then judged on data not used during that search. This makes sample-specific improvements less likely to be mistaken for robust findings. At the same time, a holdout is not a complete solution. If the holdout is consulted repeatedly and the method is revised in response, it becomes part of the search process itself. And in inference-heavy settings, holdout performance on a diagnostic criterion does not by itself validate identification, causal interpretation, or post-selection inference. The holdout should therefore be understood as a practical guardrail against hidden adaptive search, not as a full cure.